# Runyang You1    Yongqi Li1   Xinyu Lin2   Xin Zhang1 Wenjie Wang3   Wenjie Li1   Liqiang Nie4 1

### v1

@@ 5.2 Overall Performance
GRU4Rec 0.017 067 8340.0170678340.017\,067\,8340.017 067 834 0.013 454 7490.0134547490.013\,454\,7490.013 454 749 0.019 256 0180.0192560180.019\,256\,0180.019 256 018 0.014 156 5580.0141565580.014\,156\,5580.014 156 558 0.020 131 2910.0201312910.020\,131\,2910.020 131 291 0.014 363 2630.0143632630.014\,363\,2630.014 363 263 0.006 6960.0066960.006\,6960.006 696 0.003 6820.0036820.003\,6820.003 682 0.010 4170.0104170.010\,4170.010 417 0.004 1410.0041410.004\,1410.004 141 0.015 6250.0156250.015\,6250.015 625 0.005 0690.0050690.005\,0690.005 069 0.010 8640.0108640.010\,8640.010 864 0.006 9500.0069500.006\,9500.006 950 0.018 1070.0181070.018\,1070.018 107 0.009 2950.0092950.009\,2950.009 295 0.030 0740.0300740.030\,0740.030 074 0.012 3130.0123130.012\,3130.012 313
Caser 0.010 905 4080.0109054080.010\,905\,4080.010 905 408 0.014 053 9050.0140539050.014\,053\,9050.014 053 905 0.011 452 3670.0114523670.011\,452\,3670.011 452 367 0.014 912 6430.0149126430.014\,912\,6430.014 912 643 0.012 660 7920.0126607920.012\,660\,7920.012 660 792 0.015 461 320.015461320.015\,461\,320.015 461 32 0.004 464 2860.0044642860.004\,464\,2860.004 464 286 0.002 918 8690.0029188690.002\,918\,8690.002 918 869 0.006 696 4290.0066964290.006\,696\,4290.006 696 429 0.003 679 9360.0036799360.003\,679\,9360.003 679 936 0.008 928 5710.0089285710.008\,928\,5710.008 928 571 0.004 241 4690.0042414690.004\,241\,4690.004 241 469 0.012 4390.0124390.012\,4390.012 439 0.008 2630.0082630.008\,2630.008 263 0.019 0520.0190520.019\,0520.019 052 0.010 3470.0103470.010\,3470.010 347 0.027 8700.0278700.027\,8700.027 870 0.012 5740.0125740.012\,5740.012 574
SASRec 0.017 505 470.017505470.017\,505\,470.017 505 47 0.014 442 0130.0144420130.014\,442\,0130.014 442 013 0.020 131 2910.0201312910.020\,131\,2910.020 131 291 0.016 192 560.016192560.016\,192\,560.016 192 56 0.022 319 4750.0223194750.022\,319\,4750.022 319 475 0.021 006 5650.0210065650.021\,006\,5650.021 006 565 0.007 583 7320.0075837320.007\,583\,7320.007 583 732 0.010 416 6670.0104166670.010\,416\,6670.010 416 667 0.008 079 7640.0080797640.008\,079\,7640.008 079 764 0.011 904 7620.0119047620.011\,904\,7620.011 904 762 0.008 635 140.008635140.008\,635\,140.008 635 14 0.014 136 9050.0141369050.014\,136\,9050.014 136 905 0.012 9110.0129110.012\,9110.012 911 0.007 9860.0079860.007\,9860.007 986 0.020 6270.0206270.020\,6270.020 627 0.010 4510.0104510.010\,4510.010 451 0.032 5930.0325930.032\,5930.032 593 0.013 4650.0134650.013\,4650.013 465
TIGER 0.017 067 8340.0170678340.017\,067\,8340.017 067 834 0.012 779 1160.0127791160.012\,779\,1160.012 779 116 0.018 380 7440.0183807440.018\,380\,7440.018 380 744 0.013 207 3890.0132073890.013\,207\,3890.013 207 389 0.019 256 0180.0192560180.019\,256\,0180.019 256 018 0.013 436 5330.0134365330.013\,436\,5330.013 436 533 0.006 696 4290.0066964290.006\,696\,4290.006 696 429 0.004 490 7480.0044907480.004\,490\,7480.004 490 748 0.009 672 6190.0096726190.009\,672\,6190.009 672 619 0.005 468 7790.0054687790.005\,468\,7790.005 468 779 0.015 6250.0156250.015\,6250.015 625 0.006 920 2180.0069202180.006\,920\,2180.006 920 218 0.012 2820.0122820.012\,2820.012 282 0.008 4540.0084540.008\,4540.008 454 0.022 2010.0222010.022\,2010.022 201 0.011 6290.0116290.011\,6290.011 629 0.032 2780.0322780.032\,2780.032 278 0.014 1890.0141890.014\,1890.014 189
Qwen BigRec 0.005 204 944 697 462 590.005204944697462590.005\,204\,944\,697\,462\,590.005 204 944 697 462 59 0.003 322 852 180 568 223 40.00332285218056822340.003\,322\,852\,180\,568\,223\,40.003 322 852 180 568 223 4 0.011 060 507 482 108 0020.0110605074821080020.011\,060\,507\,482\,108\,0020.011 060 507 482 108 002 0.005 186 569 480 914 7830.0051865694809147830.005\,186\,569\,480\,914\,7830.005 186 569 480 914 783 0.018 867 924 528 301 8860.0188679245283018860.018\,867\,924\,528\,301\,8860.018 867 924 528 301 886 0.007 178 356 002 302 60.00717835600230260.007\,178\,356\,002\,302\,60.007 178 356 002 302 6 0.004 464 285 714 285 7140.0044642857142857140.004\,464\,285\,714\,285\,7140.004 464 285 714 285 714 0.002 468 399 825 377 9280.0024683998253779280.002\,468\,399\,825\,377\,9280.002 468 399 825 377 928 0.008 928 571 428 571 4280.0089285714285714280.008\,928\,571\,428\,571\,4280.008 928 571 428 571 428 0.003 868 539 672 861 917 50.00386853967286191750.003\,868\,539\,672\,861\,917\,50.003 868 539 672 861 917 5 0.014 136 904 761 904 7620.0141369047619047620.014\,136\,904\,761\,904\,7620.014 136 904 761 904 762 0.005 224 595 192 621 3440.0052245951926213440.005\,224\,595\,192\,621\,3440.005 224 595 192 621 344 0.000 787 640.000787640.000\,787\,640.000 787 64 0.000 388 510.000388510.000\,388\,510.000 388 51 0.001 575 20.00157520.001\,575\,20.001 575 2 0.000 641 110.000641110.000\,641\,110.000 641 11 0.012 759 9240.0127599240.012\,759\,9240.012 759 924 0.003 378 40.00337840.003\,378\,40.003 378 4
$D^{3}$ 0.004 231 770 833 333 3330.0042317708333333330.004\,231\,770\,833\,333\,3330.004 231 770 833 333 333 0.001 951 780 712 180 1950.0019517807121801950.001\,951\,780\,712\,180\,1950.001 951 780 712 180 195 0.009 440 104 166 666 6660.0094401041666666660.009\,440\,104\,166\,666\,6660.009 440 104 166 666 666 0.003 726 565 934 774 1340.0037265659347741340.003\,726\,565\,934\,774\,1340.003 726 565 934 774 134 0.019 205 729 166 666 6680.0192057291666666680.019\,205\,729\,166\,666\,6680.019 205 729 166 666 668 0.006 168 169 982 327 4740.0061681699823274740.006\,168\,169\,982\,327\,4740.006 168 169 982 327 474 0.008 184 523 809 523 810.008184523809523810.008\,184\,523\,809\,523\,810.008 184 523 809 523 81 0.005 685 264 273 880 729 50.00568526427388072950.005\,685\,264\,273\,880\,729\,50.005 685 264 273 880 729 5 0.014 136 904 761 904 7620.0141369047619047620.014\,136\,904\,761\,904\,7620.014 136 904 761 904 762 0.007 614 619 350 554 6390.0076146193505546390.007\,614\,619\,350\,554\,6390.007 614 619 350 554 639 0.025 297 619 047 619 0480.0252976190476190480.025\,297\,619\,047\,619\,0480.025 297 619 047 619 048 0.010 435 130 188 756 8430.0104351301887568430.010\,435\,130\,188\,756\,8430.010 435 130 188 756 843 0.005 356 020.005356020.005\,356\,020.005 356 02 0.002 8010.0028010.002\,8010.002 801 0.010 396 9750.0103969750.010\,396\,9750.010 396 975 0.004 394 20.00439420.004\,394\,20.004 394 2 0.019 691 240.019691240.019\,691\,240.019 691 24 0.006 707 90.00670790.006\,707\,90.006 707 9
LangPTune 0.012 6870.0126870.012\,6870.012 687 0.008 342 80.00834280.008\,342\,80.008 342 8 0.022 4460.0224460.022\,4460.022 446 0.011 4640.0114640.011\,4640.011 464 0.034 8080.0348080.034\,8080.034 808 0.014 5080.0145080.014\,5080.014 508 0.007 440 50.00744050.007\,440\,50.007 440 5 0.005 332 20.00533220.005\,332\,20.005 332 2 0.015 6250.0156250.015\,6250.015 625 0.007 990 60.00799060.007\,990\,60.007 990 6 0.020 8330.0208330.020\,8330.020 833 0.009 404 90.00940490.009\,404\,90.009 404 9 0.004 881 121 083 293 9690.0048811210832939690.004\,881\,121\,083\,293\,9690.004 881 121 083 293 969 0.002 728 30.00272830.002\,728\,30.002 728 3 0.008 815 70.00881570.008\,815\,70.008 815 7 0.003 979 20.00397920.003\,979\,20.003 979 2 0.014 0140.0140140.014\,0140.014 014 0.014 0140.0140140.014\,0140.014 014
$\text{R}^{2}\text{ec}$ 0.023 7480.0237480.023\,7480.023 748* 0.015 3850.0153850.015\,3850.015 385* 0.037 410 540.037410540.037\,410\,540.037 410 54* 0.019 7870.0197870.019\,7870.019 787* 0.061 4830.0614830.061\,4830.061 483* 0.025 8510.0258510.025\,8510.025 851* 0.051 3390.0513390.051\,3390.051 339* 0.037 1660.0371660.037\,1660.037 166* 0.064 7320.0647320.064\,7320.064 732* 0.041 3890.0413890.041\,3890.041 389* $0.0818$* 0.045 7480.0457480.045\,7480.045 748* 0.028 814 359 943 316 0130.0288143599433160130.028\,814\,359\,943\,316\,0130.028 814 359 943 316 013* 0.018 453 733 238 004 8830.0184537332380048830.018\,453\,733\,238\,004\,8830.018 453 733 238 004 883* 0.053 219 965 359 785 860.053219965359785860.053\,219\,965\,359\,785\,860.053 219 965 359 785 86* 0.026 357 850 255 487 170.026357850255487170.026\,357\,850\,255\,487\,170.026 357 850 255 487 17* 0.082 664 147 378 365 610.082664147378365610.082\,664\,147\,378\,365\,610.082 664 147 378 365 61* 0.033 710 982 335 531 660.033710982335531660.033\,710\,982\,335\,531\,660.033 710 982 335 531 66*
% Improve. 35.428 571 428 571 4135.4285714285714135.428\,571\,428\,571\,4135.428 571 428 571 41% 6.944 444 444 444 456.944444444444456.944\,444\,444\,444\,456.944 444 444 444 45% 66.964 285 714 285 7466.9642857142857466.964\,285\,714\,285\,7466.964 285 714 285 74% 22.222 222 222 222 2422.2222222222222422.222\,222\,$2 222 2$ 24% 52.605 459 057 071 9552.6054590570719552.605\,459\,057\,071\,9552.605 459 057 071 95% 23.333 333 333 333 3223.3333333333333223.333\,333\,$3 333 3$ 32% 46.571 428 571 428 55546.57142857142855546.571\,428\,571\,428\,55546.571 428 571 428 555% 58.297 872 340 425 5158.2978723404255158.297\,872\,340\,425\,5158.297 872 340 425 51% 37.953 091 684 434 96637.95309168443496637.953\,091\,684\,434\,96637.953 091 684 434 966% 51.094 890 510 948 951.094890510948951.094\,890\,510\,948\,951.094 890 510 948 9% 20.827 178 729 689 8120.8271787296898120.827\,178\,729\,689\,8120.827 178 729 689 81% 40.615 384 615 384 60640.61538461538460640.615\,384\,615\,384\,60640.615 384 615 384 606% 84.615 384 615 384 6184.6153846153846184.615\,384\,615\,384\,6184.615 384 615 384 61% 76.190 476 190 476 1876.1904761904761876.190\,476\,190\,476\,1876.190 476 190 476 18% 104.615 384 615 384 63104.61538461538463104.615\,384\,615\,384\,63104.615 384 615 384 63% 87.234 042 553 191 587.234042553191587.234\,042\,553\,191\,587.234 042 553 191 5% 92.325 581 395 348 8592.3255813953488592.325\,581\,395\,348\,8592.325 581 395 348 85% 50.446 428 571 428 58450.44642857142858450.446\,428\,571\,428\,58450.446 428 571 428 584%
Gemma BigRec 0.006 831 489 915 419 648 60.00683148991541964860.006\,831\,489\,915\,419\,648\,60.006 831 489 915 419 648 6 0.004 821 452 063 448 518 50.00482145206344851850.004\,821\,452\,063\,448\,518\,50.004 821 452 063 448 518 5 0.010 084 580 351 333 7670.0100845803513337670.010\,084\,580\,351\,333\,7670.010 084 580 351 333 767 0.005 833 862 406 063 7350.0058338624060637350.005\,833\,862\,406\,063\,7350.005 833 862 406 063 735 0.013 012 361 743 656 4730.0130123617436564730.013\,012\,361\,743\,656\,4730.013 012 361 743 656 473 0.006 600 741 721 637 5950.0066007417216375950.006\,600\,741\,721\,637\,5950.006 600 741 721 637 595 0.002 976 190 476 190 4760.0029761904761904760.002\,976\,190\,476\,190\,4760.002 976 190 476 190 476 0.002 976 190 476 190 476 50.00297619047619047650.002\,976\,190\,476\,190\,476\,50.002 976 190 476 190 476 5 0.005 208 333 333 333 3330.0052083333333333330.005\,208\,333\,333\,333\,3330.005 208 333 333 333 333 0.003 687 300 217 994 460 30.00368730021799446030.003\,687\,300\,217\,994\,460\,30.003 687 300 217 994 460 3 0.011 904 761 904 761 9040.0119047619047619040.011\,904\,761\,904\,761\,9040.011 904 761 904 761 904 0.005 312 154 716 450 0710.0053121547164500710.005\,312\,154\,716\,450\,0710.005 312 154 716 450 071 0.015 595 463 137 996 2190.0155954631379962190.015\,595\,463\,137\,996\,2190.015 595 463 137 996 219 0.010 489 160 983 015 6710.0104891609830156710.010\,489\,160\,983\,015\,6710.010 489 160 983 015 671 0.025 992 438 563 327 030.025992438563327030.025\,992\,438\,563\,327\,030.025 992 438 563 327 03 0.013 832 617 043 657 3280.0138326170436573280.013\,832\,617\,043\,657\,3280.013 832 617 043 657 328 0.043 005 671 077 504 730.043005671077504730.043\,005\,671\,077\,504\,730.043 005 671 077 504 73 0.018 162 307 633 210 0840.0181623076332100840.018\,162\,307\,633\,210\,0840.018 162 307 633 210 084
$D^{3}$ 0.007 161 458 333 333 3330.0071614583333333330.007\,161\,458\,333\,333\,3330.007 161 458 333 333 333 0.003 762 963 854 070 603 50.00376296385407060350.003\,762\,963\,854\,070\,603\,50.003 762 963 854 070 603 5 0.020 182 291 666 666 6680.0201822916666666680.020\,182\,291\,666\,666\,6680.020 182 291 666 666 668 0.007 995 190 196 701 8170.0079951901967018170.007\,995\,190\,196\,701\,8170.007 995 190 196 701 817 0.033 854 166 666 666 6640.0338541666666666640.033\,854\,166\,666\,666\,6640.033 854 166 666 666 664 0.011 421 699 038 261 1520.0114216990382611520.011\,421\,699\,038\,261\,1520.011 421 699 038 261 152 0.021 577 380 952 380 9520.0215773809523809520.021\,577\,380\,952\,380\,9520.021 577 380 952 380 952 0.012 923 434 799 794 6520.0129234347997946520.012\,923\,434\,799\,794\,6520.012 923 434 799 794 652 0.032 738 095 238 095 240.032738095238095240.032\,738\,095\,238\,095\,240.032 738 095 238 095 24 0.016 384 061 220 719 0350.0163840612207190350.016\,384\,061\,220\,719\,0350.016 384 061 220 719 035 0.044 642 857 142 857 1440.0446428571428571440.044\,642\,857\,142\,857\,1440.044 642 857 142 857 144 0.019 403 363 128 620 0520.0194033631286200520.019\,403\,363\,128\,620\,0520.019 403 363 128 620 052 0.011 657 214 870 825 4580.0116572148708254580.011\,657\,214\,870\,825\,4580.011 657 214 870 825 458 0.006 789 634 641 773 0710.0067896346417730710.006\,789\,634\,641\,773\,0710.006 789 634 641 773 071 0.020 951 480 781 348 4560.0209514807813484560.020\,951\,480\,781\,348\,4560.020 951 480 781 348 456 0.014 083 412 140 940 2040.0140834121409402040.014\,083\,412\,140\,940\,2040.014 083 412 140 940 204 0.037 807 183 364 839 320.037807183364839320.037\,807\,183\,364\,839\,320.037 807 183 364 839 32 0.022 433 535 336 669 4640.0224335353366694640.022\,433\,535\,336\,669\,4640.022 433 535 336 669 464
LangPTune 0.013 012 3620.0130123620.013\,012\,3620.013 012 362 0.007 850 0180.0078500180.007\,850\,0180.007 850 018 0.022 121 014 964 2160.0221210149642160.022\,121\,014\,964\,2160.022 121 014 964 216 0.010 718 4340.0107184340.010\,718\,4340.010 718 434 0.040 338 321 405 3350.0403383214053350.040\,338\,321\,405\,3350.040 338 321 405 335 0.015 231 8650.0152318650.015\,231\,8650.015 231 865 0.034 970 238 095 238 0960.0349702380952380960.034\,970\,238\,095\,238\,0960.034 970 238 095 238 096 0.023 467 457 188 027 2470.0234674571880272470.023\,467\,457\,188\,027\,2470.023 467 457 188 027 247 0.046 8750.0468750.046\,8750.046 875 0.027 365 138 310 761 6770.0273651383107616770.027\,365\,138\,310\,761\,6770.027 365 138 310 761 677 0.067 708 333 333 333 330.067708333333333330.067\,708\,333\,333\,333\,330.067 708 333 333 333 33 0.032 544 824 028 653 760.032544824028653760.032\,544\,824\,028\,653\,760.032 544 824 028 653 76 0.006 770 60.00677060.006\,770\,60.006 770 6 0.005 332 20.00533220.005\,332\,20.005 332 2 0.011 9670.0119670.011\,9670.011 967 0.005 852 70.00585270.005\,852\,70.005 852 7 0.019 5240.0195240.019\,5240.019 524 0.009 404 90.00940490.009\,404\,90.009 404 9
$\text{R}^{2}\text{ec}$ 0.026 350.026350.026\,350.026 35* 0.016 0810.0160810.016\,0810.016 081* 0.039 6880.0396880.039\,6880.039 688* 0.020 3050.0203050.020\,3050.020 305* 0.061 4830.0614830.061\,4830.061 483* 0.025 7280.0257280.025\,7280.025 728* 0.057 2920.0572920.057\,2920.057 292* 0.039 7510.0397510.039\,7510.039 751* 0.080 3570.0803570.080\,3570.080 357* 0.047 1540.0471540.047\,1540.047 154* 0.104 170.104170.104\,170.104 17* 0.052 6720.0526720.052\,6720.052 672* 0.032 593 292 394 898 440.032593292394898440.032\,593\,292\,394\,898\,440.032 593 292 394 898 44* 0.020 513 333 770 876 2880.0205133337708762880.020\,513\,333\,770\,876\,2880.020 513 333 770 876 288* 0.053 062 509 840 969 9260.0530625098409699260.053\,062\,509\,840\,969\,9260.053 062 509 840 969 926* 0.027 052 420 548 539 7840.0270524205485397840.027\,052\,420\,548\,539\,7840.027 052 420 548 539 784* 0.083 451 424 972 445 290.083451424972445290.083\,451\,424\,972\,445\,290.083 451 424 972 445 29* 0.034 727 909 699 741 6760.0347279096997416760.034\,727\,909\,699\,741\,6760.034 727 909 699 741 676*
[…]

### v3

@@ 4.2 Overall Performance
GRU4Rec 0.017 067 8340.017\,067\,834 0.013 454 7490.013\,454\,749 0.019 256 0180.019\,256\,018 0.014 156 5580.014\,156\,558 0.020 131 2910.020\,131\,291 0.014 363 2630.014\,363\,263 0.006 6960.006\,696 0.003 6820.003\,682 0.010 4170.010\,417 0.004 1410.004\,141 0.015 6250.015\,625 0.005 0690.005\,069 0.010 8640.010\,864 0.006 9500.006\,950 0.018 1070.018\,107 0.009 2950.009\,295 0.030 0740.030\,074 0.012 3130.012\,313
Caser 0.010 905 4080.010\,905\,408 0.014 053 9050.014\,053\,905 0.011 452 3670.011\,452\,367 0.014 912 6430.014\,912\,643 0.012 660 7920.012\,660\,792 0.015 461 320.015\,461\,32 0.004 464 2860.004\,464\,286 0.002 918 8690.002\,918\,869 0.006 696 4290.006\,696\,429 0.003 679 9360.003\,679\,936 0.008 928 5710.008\,928\,571 0.004 241 4690.004\,241\,469 0.012 4390.012\,439 0.008 2630.008\,263 0.019 0520.019\,052 0.010 3470.010\,347 0.027 8700.027\,870 0.012 5740.012\,574
SASRec 0.017 505 470.017\,505\,47 0.014 442 0130.014\,442\,013 0.020 131 2910.020\,131\,291 0.016 192 560.016\,192\,56 0.022 319 4750.022\,319\,475 0.021 006 5650.021\,006\,565 0.007 583 7320.007\,583\,732 0.010 416 6670.010\,416\,667 0.008 079 7640.008\,079\,764 0.011 904 7620.011\,904\,762 0.008 635 140.008\,635\,14 0.014 136 9050.014\,136\,905 0.012 9110.012\,911 0.007 9860.007\,986 0.020 6270.020\,627 0.010 4510.010\,451 0.032 5930.032\,593 0.013 4650.013\,465
TIGER 0.017 067 8340.017\,067\,834 0.012 779 1160.012\,779\,116 0.018 380 7440.018\,380\,744 0.013 207 3890.013\,207\,389 0.019 256 0180.019\,256\,018 0.013 436 5330.013\,436\,533 0.006 696 4290.006\,696\,429 0.004 490 7480.004\,490\,748 0.009 672 6190.009\,672\,619 0.005 468 7790.005\,468\,779 0.015 6250.015\,625 0.006 920 2180.006\,920\,218 0.012 2820.012\,282 0.008 4540.008\,454 0.022 2010.022\,201 0.011 6290.011\,629 0.032 2780.032\,278 0.014 1890.014\,189
Qwen
BigRec 0.005 204 944 697 462 590.005\,204\,944\,697\,462\,59 0.003 322 852 180 568 223 40.003\,322\,852\,180\,568\,223\,4 0.011 060 507 482 108 0020.011\,060\,507\,482\,108\,002 0.005 186 569 480 914 7830.005\,186\,569\,480\,914\,783 0.018 867 924 528 301 8860.018\,867\,924\,528\,301\,886 0.007 178 356 002 302 60.007\,178\,356\,002\,302\,6 0.004 464 285 714 285 7140.004\,464\,285\,714\,285\,714 0.002 468 399 825 377 9280.002\,468\,399\,825\,377\,928 0.008 928 571 428 571 4280.008\,928\,571\,428\,571\,428 0.003 868 539 672 861 917 50.003\,868\,539\,672\,861\,917\,5 0.014 136 904 761 904 7620.014\,136\,904\,761\,904\,762 0.005 224 595 192 621 3440.005\,224\,595\,192\,621\,344 0.000 787 640.000\,787\,64 0.000 388 510.000\,388\,51 0.001 575 20.001\,575\,2 0.000 641 110.000\,641\,11 0.012 759 9240.012\,759\,924 0.003 378 40.003\,378\,4
$D^{3}$ 0.004 231 770 833 333 3330.004\,231\,770\,833\,333\,333 0.001 951 780 712 180 1950.001\,951\,780\,712\,180\,195 0.009 440 104 166 666 6660.009\,440\,104\,166\,666\,666 0.003 726 565 934 774 1340.003\,726\,565\,934\,774\,134 0.019 205 729 166 666 6680.019\,205\,729\,166\,666\,668 0.006 168 169 982 327 4740.006\,168\,169\,982\,327\,474 0.008 184 523 809 523 810.008\,184\,523\,809\,523\,81 0.005 685 264 273 880 729 50.005\,685\,264\,273\,880\,729\,5 0.014 136 904 761 904 7620.014\,136\,904\,761\,904\,762 0.007 614 619 350 554 6390.007\,614\,619\,350\,554\,639 0.025 297 619 047 619 0480.025\,297\,619\,047\,619\,048 0.010 435 130 188 756 8430.010\,435\,130\,188\,756\,843 0.005 356 020.005\,356\,02 0.002 8010.002\,801 0.010 396 9750.010\,396\,975 0.004 394 20.004\,394\,2 0.019 691 240.019\,691\,24 0.006 707 90.006\,707\,9
LangPTune 0.012 6870.012\,687 0.008 342 80.008\,342\,8 0.022 4460.022\,446 0.011 4640.011\,464 0.034 8080.034\,808 0.014 5080.014\,508 0.007 440 50.007\,440\,5 0.005 332 20.005\,332\,2 0.015 6250.015\,625 0.007 990 60.007\,990\,6 0.020 8330.020\,833 0.009 404 90.009\,404\,9 0.004 881 121 083 293 9690.004\,881\,121\,083\,293\,969 0.002 728 30.002\,728\,3 0.008 815 70.008\,815\,7 0.003 979 20.003\,979\,2 0.014 0140.014\,014 0.014 0140.014\,014
$\text{R}^{2}\text{ec}$ 0.023 7480.023\,748* 0.015 3850.015\,385* 0.037 410 540.037\,410\,54* 0.019 7870.019\,787* 0.061 4830.061\,483* 0.025 8510.025\,851* 0.051 3390.051\,339* 0.037 1660.037\,166* 0.064 7320.064\,732* 0.041 3890.041\,389* 0.08180.0818* 0.045 7480.045\,748* 0.028 814 359 943 316 0130.028\,814\,359\,943\,316\,013* 0.018 453 733 238 004 8830.018\,453\,733\,238\,004\,883* 0.053 219 965 359 785 860.053\,219\,965\,359\,785\,86* 0.026 357 850 255 487 170.026\,357\,850\,255\,487\,17* 0.082 664 147 378 365 610.082\,664\,147\,378\,365\,61* 0.033 710 982 335 531 660.033\,710\,982\,335\,531\,66*
% Improve. 35.428 571 428 571 4135.428\,571\,428\,571\,41% 6.944 444 444 444 456.944\,444\,444\,444\,45% 66.964 285 714 285 7466.964\,285\,714\,285\,74% 22.222 222 222 222 2422.222\,222\,222\,222\,24% 52.605 459 057 071 9552.605\,459\,057\,071\,95% 23.333 333 333 333 3223.333\,333\,333\,333\,32% 46.571 428 571 428 55546.571\,428\,571\,428\,555% 58.297 872 340 425 5158.297\,872\,340\,425\,51% 37.953 091 684 434 96637.953\,091\,684\,434\,966% 51.094 890 510 948 951.094\,890\,510\,948\,9% 20.827 178 729 689 8120.827\,178\,729\,689\,81% 40.615 384 615 384 60640.615\,384\,615\,384\,606% 42.361 111 142.361\,111\,1% 34.054 054 054 054 034.054\,054\,054\,054\,0% 51.127 819 551.127\,819\,5% 41.287 878 787 878 741.287\,878\,787\,878\,7% 31.559 854 897 231.559\,854\,897\,2% 33.531 157 270 033.531\,157\,270\,0%
Gemma
BigRec 0.006 831 489 915 419 648 60.006\,831\,489\,915\,419\,648\,6 0.004 821 452 063 448 518 50.004\,821\,452\,063\,448\,518\,5 0.010 084 580 351 333 7670.010\,084\,580\,351\,333\,767 0.005 833 862 406 063 7350.005\,833\,862\,406\,063\,735 0.013 012 361 743 656 4730.013\,012\,361\,743\,656\,473 0.006 600 741 721 637 5950.006\,600\,741\,721\,637\,595 0.002 976 190 476 190 4760.002\,976\,190\,476\,190\,476 0.002 976 190 476 190 476 50.002\,976\,190\,476\,190\,476\,5 0.005 208 333 333 333 3330.005\,208\,333\,333\,333\,333 0.003 687 300 217 994 460 30.003\,687\,300\,217\,994\,460\,3 0.011 904 761 904 761 9040.011\,904\,761\,904\,761\,904 0.005 312 154 716 450 0710.005\,312\,154\,716\,450\,071 0.015 595 463 137 996 2190.015\,595\,463\,137\,996\,219 0.010 489 160 983 015 6710.010\,489\,160\,983\,015\,671 0.025 992 438 563 327 030.025\,992\,438\,563\,327\,03 0.013 832 617 043 657 3280.013\,832\,617\,043\,657\,328 0.043 005 671 077 504 730.043\,005\,671\,077\,504\,73 0.018 162 307 633 210 0840.018\,162\,307\,633\,210\,084
$D^{3}$ 0.007 161 458 333 333 3330.007\,161\,458\,333\,333\,333 0.003 762 963 854 070 603 50.003\,762\,963\,854\,070\,603\,5 0.020 182 291 666 666 6680.020\,182\,291\,666\,666\,668 0.007 995 190 196 701 8170.007\,995\,190\,196\,701\,817 0.033 854 166 666 666 6640.033\,854\,166\,666\,666\,664 0.011 421 699 038 261 1520.011\,421\,699\,038\,261\,152 0.021 577 380 952 380 9520.021\,577\,380\,952\,380\,952 0.012 923 434 799 794 6520.012\,923\,434\,799\,794\,652 0.032 738 095 238 095 240.032\,738\,095\,238\,095\,24 0.016 384 061 220 719 0350.016\,384\,061\,220\,719\,035 0.044 642 857 142 857 1440.044\,642\,857\,142\,857\,144 0.019 403 363 128 620 0520.019\,403\,363\,128\,620\,052 0.011 657 214 870 825 4580.011\,657\,214\,870\,825\,458 0.006 789 634 641 773 0710.006\,789\,634\,641\,773\,071 0.020 951 480 781 348 4560.020\,951\,480\,781\,348\,456 0.014 083 412 140 940 2040.014\,083\,412\,140\,940\,204 0.047 807 183 364 839 320.047\,807\,183\,364\,839\,32 0.022 433 535 336 669 4640.022\,433\,535\,336\,669\,464
SDPO* $0.0066$ $0.0034$ $0.0098$ $0.0054$ $0.0144$ $0.0071$ $0.0022$ $0.0018$ $0.0037$ $0.0025$ $0.0162$ $0.0094$ $0.0166$ $0.0122$ $0.0298$ $0.0155$ $0.0466$ $0.0222$
Llara* 0.007 807 417 046 193 8840.007\,807\,417\,046\,193\,884 0.005 536 802 492 208 3570.005\,536\,802\,492\,208\,357 0.013 662 979 830 839 2970.013\,662\,979\,830\,839\,297 0.007 443 086 853 763 2470.007\,443\,086\,853\,763\,247 0.015 940 143 135 979 180.015\,940\,143\,135\,979\,18 0.008 002 939 751 013 4260.008\,002\,939\,751\,013\,426 $0.0097$ $0.0039$ $0.0127$ $0.0049$ $0.0202$ $0.0152$ 0.027 455 655 60.027\,455\,655\,6 0.017 257 1960.017\,257\,196 0.042 798 490.042\,798\,49 0.022 280 000.022\,280\,$0$ 0.067 722 456 60.067\,722\,456\,6 0.029 850 110.029\,850\,11
SPRec $0.0070$ $0.0033$ $0.0111$ $0.0062$ $0.0142$ $0.0077$ $0.0029$ $0.0022$ $0.0037$ $0.0025$ $0.0124$ $0.0063$ $0.0152$ $0.0113$ $0.0244$ $0.0133$ $0.0566$ $0.0211$
LangPTune 0.013 012 3620.013\,012\,362 0.007 850 0180.007\,850\,018 0.022 121 014 964 2160.022\,121\,014\,964\,216 0.010 718 4340.010\,718\,434 0.040 338 321 405 3350.040\,338\,321\,405\,335 0.015 231 8650.015\,231\,865 0.034 970 238 095 238 0960.034\,970\,238\,095\,238\,096 0.023 467 457 188 027 2470.023\,467\,457\,188\,027\,247 0.046 8750.046\,875 0.027 365 138 310 761 6770.027\,365\,138\,310\,761\,677 0.067 708 333 333 333 330.067\,708\,333\,333\,333\,33 0.032 544 824 028 653 760.032\,544\,824\,028\,653\,76 0.006 770 60.006\,770\,6 0.005 332 20.005\,332\,2 0.011 9670.011\,967 0.005 852 70.005\,852\,7 0.019 5240.019\,524 0.009 404 90.009\,404\,9
$\text{R}^{2}\text{ec}$ 0.026 350.026\,35* 0.016 0810.016\,081* 0.039 6880.039\,688* 0.020 3050.020\,305* 0.061 4830.061\,483* 0.025 7280.025\,728* 0.057 2920.057\,292* 0.039 7510.039\,751* 0.080 3570.080\,357* 0.047 1540.047\,154* 0.104 170.104\,17* 0.052 6720.052\,672* 0.032 593 292 394 898 440.032\,593\,292\,394\,898\,44* 0.020 513 333 770 876 2880.020\,513\,333\,770\,876\,288* 0.053 062 509 840 969 9260.053\,062\,509\,840\,969\,926* 0.027 052 420 548 539 7840.027\,052\,420\,548\,539\,784* 0.083 451 424 972 445 290.083\,451\,424\,972\,445\,29* 0.034 727 909 699 741 6760.034\,727\,909\,699\,741\,676*
[…]